A body covered a distance of L m along a curved path of a quarter circle.The ratio of distance to displacement will be?? 

Grade:11

2 Answers

Kuldip Jha
7 years ago
total distance travelled=L
for displacement (pi)r/2= L
so r = 2L/pi
dispacement=shortest distance between initial and final points=r=2\sqrt{2}L/pi
therefore ratio=pi/2\sqrt{2}
bulbul das
6 years ago
total distance travelled = l
for displacement = pi r/2=l
so r = 2l/pi
displacement = shortest distance between initial and final points = r = 2root2l/pi
therefore ratio = pi/2root2